OpenStudy (anonymous):

The coordinates of quadrilateral FGHJ are F(-5, -3), G(-3, 3), H(3, 5), and J(7, -1). Find the midpoint of FG

jimthompson5910 (jim_thompson5910):

Use the formula C = pi*d where d = diameter

OpenStudy (anonymous):

\[c=\pi*d\]

OpenStudy (anonymous):

yes @jim_thompson5910 got it right

OpenStudy (anonymous):

how can it be c c=pi*d 4pi = pi*d pi gets canceled so d=4 answer is option D

OpenStudy (anonymous):

Oh Ok I multiplied the wrong numbers

OpenStudy (anonymous):

the formula they gave you is c=2pi*r but 2*radius = diameter right? don't get confused btwn basics <3